OpenStudy (anonymous):

solve for x: 8^x=5^(2x-1)

OpenStudy (anonymous):

1. take natural log of both sides

OpenStudy (anonymous):

How would I take the natural log of 5^(2x-1)

OpenStudy (anonymous):

ln(5^(2x-1)) = (2x-1)ln(5)

OpenStudy (anonymous):

using law of logarithms

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

x(ln(8))=2x-1(ln(5))

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

Next, I would divide by 2x-1?

OpenStudy (anonymous):

And ln(8) on both sides?

OpenStudy (anonymous):

either way you do it, I am sure you will get the right answer.

OpenStudy (anonymous):

I got ln(5/8)=x/2x-1

OpenStudy (anonymous):

x/(2x-1)

OpenStudy (anonymous):

What would I do from there?

OpenStudy (anonymous):

try solving for x using proportion rules

OpenStudy (anonymous):

try it this way will be easier: \[x(\frac{ \ln8 }{ \ln5 })-2x+1=0\]

OpenStudy (anonymous):

move everything to one side